### v1

# OmniFC: Rethinking Federated Clustering via Lossless and Secure Distance Reconstruction

Abstract
Federated clustering (FC) aims to discover global cluster structures across decentralized clients without sharing raw data, making privacy preservation a fundamental requirement. There are two critical challenges: (1) privacy leakage during collaboration, and (2) robustness degradation due to aggregation of proxy information from non-independent and identically distributed (Non-IID) local data, leading to inaccurate or inconsistent global clustering. Existing solutions typically rely on model-specific local proxies, which are sensitive to data heterogeneity and inherit inductive biases from their centralized counterparts, thus limiting robustness and generality. We propose Omni Federated Clustering (OmniFC), a unified and model-agnostic framework. Leveraging Lagrange coded computing, our method enables clients to share only encoded data, allowing exact reconstruction of the global distance matrix—a fundamental representation of sample relationships—without leaking private information, even under client collusion. This construction is naturally resilient to Non-IID data distributions. This approach decouples FC from model-specific proxies, providing a unified extension mechanism applicable to diverse centralized clustering methods. Theoretical analysis confirms both reconstruction fidelity and privacy guarantees, while comprehensive experiments demonstrate OmniFC’s superior robustness, effectiveness, and generality across various benchmarks compared to state-of-the-art methods. Code will be released.
1 Introduction
Traditional clustering methods presuppose centralized access to the entire dataset, enabling the construction of global structures such as cluster centroids or kernel matrices. However, in federated settings characterized by data fragmentation across clients and privacy constraints, this assumption breaks down, precluding direct application.
To overcome this, federated clustering (FC) dennis2021heterogeneity has emerged, where clients collaboratively group data without sharing raw samples. There are two fundamental challenges: (1) privacy leakage during collaboration, and (2) robustness degradation under non-independent and identically distributed (Non-IID) data. Existing FC methods approximate global structures by aggregating model-specific local proxies: federated k-means (KM) and fuzzy c-means (FCM) aggregate local cluster centroids dennis2021heterogeneity ; stallmann2022towards ; pan2023machine ; xu2024jigsaw ; yang2024greedy , federated spectral clustering (SC) qiao2023federated reconstructs the global kernel matrix from local low-rank factors, and federated non-negative matrix factorization (NMF) wang2022federated aggregates local basis matrices. These proxies, however, are computed from biased client-specific datasets, fail to reliably capture global structures, leading to degraded robustness and performance (Fig. 1). Moreover, such methods are tightly bound to specific centralized clustering methods, inheriting restrictive inductive biases—e.g., data compactness in KM macqueen1967some and FCM bezdek1984fcm , data connectivity in SC ng2001spectral , and low-rank representation in NMF lee2000algorithms —thereby confining their performance to assumption-compliant data and limiting their generality.
Figure 1: Robustness to heterogeneity. We employ COIL-100 and 100 clients to compare the proposed OmniFC with the federated extensions of centralized clustering methods dennis2021heterogeneity ; pan2023machine ; stallmann2022towards ; qiao2023federated ; wang2022federated . Compared to existing one-to-one extensions, OmniFC not only unifies the extension of centralized clustering methods but also achieves superior robustness and effectiveness.
This work addresses both limitations through a unifying perspective: reconstructing the global pairwise distance matrix, which offers a model-agnostic and fundamental representation of sample relationships, naturally resilient to the Non-IID problem. The key challenge, however, lies in securely computing this matrix without exposing private data. To this end, we propose Omni Federated Clustering (OmniFC), a novel framework that facilitates a unified extension from centralized clustering to FC through lossless and secure distance reconstruction. OmniFC comprises three main steps: local Lagrange-encoded sharing, global distance reconstruction, and cluster assignment. Each client initially encrypts its local data using Lagrange coded computing yu2019lagrange , shares the encoded data with peers for pairwise distance computation, and subsequently transmits the resulting distances to the central server for constructing the global distance matrix. Finally, the global distance matrix can serve as input to centralized clustering methods for performing cluster assignment. Fig. 1 demonstrates the superiority of OmniFC. With respect to distance reconstruction, the proposed OmniFC exhibits two salient features: 1) Efficacy. Both theoretical and empirical analyses consistently demonstrate the capability for lossless reconstruction and robustness to the Non-IID problem. Benefiting from this, the proposed OmniFC achieves lossless federated extensions for pairwise-distance-dependent methods (e.g., SC) and enhances federated extensions for methods (e.g., KM) without explicit dependence on pairwise distances. 2) Security. Theoretical analysis demonstrates that the privacy of local data is preserved during data sharing, as the encoded data prevents the inference of private information even under client collusion. In summary, our contributions are threefold:
1) We propose OmniFC, a novel framework that facilitates a unified extension from centralized clustering to FC through lossless and secure distance reconstruction.
2) We establish theoretical assurances regarding the efficacy and security of distance reconstruction.
3) Experimental results show that our OmniFC outperforms SOTA methods on various benchmarks.
2 Related Work
Centralized Clustering.
Traditional centralized clustering aggregates client-held local data on a central server for grouping, with methods making different assumptions—such as compactness macqueen1967some ; ikotun2023k , connectivity ng2001spectral ; ding2024survey , density ester1996density ; xu2024scalable , hierarchy gowda1978agglomerative ; laber2024cohesion , and low-rank representation lee2000algorithms ; li2024tensorized of the data distribution—to adapt to diverse datasets. However, these methods may become inapplicable due to privacy constraints that prevent the centralization of client data.
Federated Clustering (FC).
Unlike centralized clustering, which requires collecting raw client data for model training, FC collects local proxies instead, thus strengthening user privacy protection. To handle this, several recent works have shifted from sharing local private data to exchanging local cluster centroids dennis2021heterogeneity ; stallmann2022towards ; xu2024jigsaw ; yang2024greedy , local basis matrices wang2022federated or synthetic data yan2024sda . Although these methods show promise, these methods either suffer from performance degradation caused by the Non-IID problem or achieve gains at the expense of privacy yan2024sda .
Secure FC.
Secure FC leverages advanced privacy-preserving techniques—including differential privacy dwork2014algorithmic , machine unlearning cao2015towards , and Lagrange coded computing yu2019lagrange —to concurrently improve clustering efficacy and fortify data confidentiality. Existing methods typically focus on the effective and secure construction of either global cluster centroids for k-means li2022secure ; pan2023machine ; wang2024one or a global kernel matrix for spectral clustering qiao2023federated . Although promising, these methods remain limited by the Non-IID problem or fail to offer a model-agnostic solution. Moreover, they inherently retain assumptions—such as data compactness macqueen1967some ; ikotun2023k and connectivity ng2001spectral ; ding2024survey —from their centralized counterparts, limiting their effectiveness to compliant datasets and thereby reducing their practical applicability. To handle these, we introduce Omni Federated Clustering (OmniFC), a unified and model-agnostic framework. Leveraging Lagrange coded computing, our method enables clients to share only encoded data, allowing exact reconstruction of the global distance matrix—a fundamental representation of sample relationships—without leaking private information, even under client collusion. This construction is naturally resilient to the Non-IID data distributions. This approach decouples FC from model-specific proxies, providing a unified extension mechanism that can be applied to various centralized clustering methods to accommodate diverse datasets.
3 Omni Federated Clustering (OmniFC)
This section begins with an overview of the problem definition and the OmniFC framework, followed by a detailed description of OmniFC, and concludes with its privacy analysis.
3.1 Overview
Problem Definition.
[…]
Figure 2: An overview of the proposed OmniFC. The architecture comprises three main steps: 1) Local Lagrange-Encoded Sharing. Each client $j$ $(j\in[m])$ encodes its private data using Lagrange polynomial interpolation and distributes the encoded data to all peers, enabling each client to construct a global encoded dataset while preserving data privacy. 2) Global Distance Reconstruction. Each client $j$ computes pairwise distances within its global encoded dataset and transmits the results to the central server, which leverages them to reconstruct the global distance matrix. 3) Cluster Assignment. A centralized clustering method (e.g., k-means) is applied to the global distance matrix to produce the final clustering result $\pi^{*}$.
Framework Overview.
As shown in Fig. 2, OmniFC comprises three main steps: local Lagrange-encoded sharing, global distance reconstruction, and cluster assignment. Each client $j$ $(j\in[m])$ initially encrypts its local data using Lagrange coded computing (LCC) yu2019lagrange , shares the encoded data with peers for pairwise distance computation, and subsequently transmits the resulting distances to the central server for constructing the global distance matrix. Finally, the global distance matrix can serve as input to centralized clustering methods for performing cluster assignment.
3.2 OmniFC
Local Lagrange-encoded Sharing.
First, each sample $\boldsymbol{x}_{i}\in\boldsymbol{X}$ $(i\in[n])$—regardless of the client to which it is distributed to—is independently transformed from the real domain $\mathbb{R}^{d}$ to the finite field $\mathbb{F}_{p}^{d}$ to ensure numerical stability in secure computation yu2019lagrange , with $p$ denoting a prime. The transformation is defined as:
$\boldsymbol{\tilde{x}}_{i}=round(2^{q}\cdot\boldsymbol{x}_{i})+p\cdot\frac{|sign(\boldsymbol{x}_{i})|-sign(\boldsymbol{x}_{i})}{2},$ (1)
where $q\in\mathbb{Z}$ regulates the quantization loss. $round(\cdot)$ and $sign(\cdot)$ represent element-wise rounding and sign functions, respectively. Rounding discretizes continuous values to ensure finite field compatibility, while the sign function facilitates correct mapping of negative values shao2022dres . We denote the transformed form of $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ as $\boldsymbol{\tilde{X}}\in\mathbb{F}_{p}^{n\times d}$.
Then, each sample $\boldsymbol{\tilde{x}}_{i}\in\boldsymbol{\tilde{X}}$ $(i\in[n])$ is independently encoded via Lagrange polynomial interpolation by the client (Fig. 3), enabling secret sharing among clients. Specifically, $\boldsymbol{\tilde{x}}_{i}\in\mathbb{F}_{p}^{d}$ is partitioned into $l$ segments $\{\boldsymbol{s}_{i,o}\}_{o=1}^{l}$, i.e., $\boldsymbol{\tilde{x}}_{i}=[\boldsymbol{s}_{i,1}^{T},\boldsymbol{s}_{i,2}^{T},\cdots,\boldsymbol{s}_{i,l}^{T}]^{T}$, and combined with $t$ random noises to construct a polynomial that serves to encode $\boldsymbol{\tilde{x}}_{i}$. The noises are introduced to ensure privacy protection against potential client collusion yu2019lagrange . Assuming that $d$ is divisible by $l$, the client holds data segments $\boldsymbol{s}_{i,o}\in\mathbb{F}_{p}^{\frac{d}{l}}$ $(o\in[l])$, and samples $t$ additional noises $\boldsymbol{s}_{i,l+o}$ $(o\in[t])$ uniformly from $\mathbb{F}_{p}^{\frac{d}{l}}$. Based on the segments $\{\boldsymbol{s}_{i,o}\}_{o=1}^{l+t}$, the Lagrange interpolation polynomial $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}:\mathbb{F}_{p}\rightarrow\mathbb{F}_{p}^{\frac{d}{l}}$ of degree $l+t-1$ can be constructed as follows:
$\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\alpha)=\sum_{o=1}^{l+t}\boldsymbol{s}_{i,o}\cdot\prod_{{o^{\prime}}\neq o}\frac{\alpha-\alpha_{o^{\prime}}}{\alpha_{o}-\alpha_{o^{\prime}}},$ (2)
where $\{{\alpha}_{o}\}_{o=1}^{l+t}$ denotes a set of $l+t$ distinct hyperparameters from $\mathbb{F}_{p}$, pre-specified through agreement among all clients and the central server. Particularly, each data segment $\boldsymbol{s}_{i,o}$ $(o\in[l])$ can be recovered by setting $\alpha={\alpha}_{o}$, i.e., $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\alpha_{o})=s_{i,o}$. Beyond the $\{{\alpha}_{o}\}_{o=1}^{l+t}$ employed in constructing the polynomial $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}$, all clients and the central server also pre-select $m$ distinct public hyperparameters $\{{\beta}_{j}\}_{j=1}^{m}$ for encoding, where ${\beta}_{j}\in\mathbb{F}_{p}$ and $\{{\alpha}_{o}\}_{o=1}^{l+t}\cap\{{\beta}_{j}\}_{j=1}^{m}=\emptyset$. Based on $\{{\beta}_{j}\}_{j=1}^{m}$, the client encodes its local data $\boldsymbol{\tilde{x}}_{i}$ into $m$ distinct representations $\{\boldsymbol{z}_{i,j}\}_{j=1}^{m}$ for secret sharing, with each representation
$\boldsymbol{z}_{i,j}=\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\beta_{j})$ (3)
delivered to the $j$-th client.
As these operations are defined per sample, they are universally applicable to local data across all clients. Hence, each client $j$ $(j\in[m])$ will possess a global encoded dataset $\boldsymbol{Z}_{j}\in\mathbb{F}_{p}^{n\times\frac{d}{l}}$ corresponding to $\boldsymbol{\tilde{X}}\in\mathbb{F}_{p}^{n\times d}$, where $\boldsymbol{Z}_{j}=[\boldsymbol{z}_{1,j},\boldsymbol{z}_{2,j},\cdots,\boldsymbol{z}_{n,j}]^{T}=[\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{1}}(\beta_{j}),\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{2}}(\beta_{j}),...,\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{n}}(\beta_{j})]^{T}$.
Figure 3: An illustration of the Lagrange encoding. Each sample $\boldsymbol{\tilde{x}}_{i}$ $(i\in[n])$ is initially divided into $l$ segments $\{\boldsymbol{s}_{i,o}\}_{o=1}^{l}$. Incorporating $t$ additional noises $\{\boldsymbol{s}_{i,l+o}\}_{o=1}^{t}$, Lagrange interpolation is then conducted as per Equation (2) to yield $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\alpha)$. Subsequently, the encoded representations $\{\boldsymbol{z}_{i,j}\}_{j=1}^{m}$ of $\boldsymbol{\tilde{x}}_{i}$ are computed according to Equation (3).
Global Distance Reconstruction.
For each client $j$ $(j\in[m])$, pairwise distances between all encoded representations $\boldsymbol{z}_{i,j}$ and $\boldsymbol{z}_{i^{\prime},j}$ in $\boldsymbol{Z}_{j}$ $(i,i^{\prime}\in[n])$ are calculated and subsequently sent to the central server for constructing the global distance matrix. Specifically, the pairwise distance between $\boldsymbol{z}_{i,j}$ and $\boldsymbol{z}_{i^{\prime},j}$ can be calculated as:
${dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j})=\left\|\boldsymbol{z}_{i,j}-\boldsymbol{z}_{i^{\prime},j}\right\|_{2}^{2},$ (4)
Based on the $m$ distances $\{{dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j})\}_{j=1}^{m}$ provided by the clients, the server can accurately recover the pairwise distance between the corresponding samples $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{\tilde{x}}_{i^{\prime}}$, as demonstrated in Theorem 1.
Theorem 1.
Let $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\beta):\mathbb{F}_{p}\rightarrow\mathbb{F}_{p}$ denote the Lagrange interpolation polynomial interpolated from the set $\{(\beta_{j},{dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j})\}_{j=1}^{m}$, where $\boldsymbol{z}_{i,j}$ and $\boldsymbol{z}_{i^{\prime},j}$ denote the encoded representations of arbitrary samples $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{\tilde{x}}_{i^{\prime}}$ distributed among clients. When $m\geq 2l+2t-1$, the distance $\text{dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})$ can be precisely recovered:
${dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})=\sum_{o=1}^{l}\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\alpha_{o}),$ (5)
irrespective of how data is distributed among clients.
Remark 1.
The condition $m\geq 2l+2t-1$ imposes minimal practical constraint, given that $m$ is predefined by the system while $l$ and $t$ are tunable hyperparameters. This flexibility allows the condition to be met easily, ensuring the theorem’s practical applicability and highlighting its relevance to real-world implementations.
Then, by converting ${dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})$ from the finite field $\mathbb{F}_{p}$ back to the real domain $\mathbb{R}$, the server recovers:
$\left.{dis}(\boldsymbol{x}_{i},\boldsymbol{x}_{i^{\prime}})=\left\{\begin{array}[]{lll}\frac{1}{2^{q}}\cdot{dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})&\mathrm{if}&0\leq{dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})<\frac{p-1}{2}\\[10.0pt] \frac{1}{2^{q}}\cdot({dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})-p)&\mathrm{if}&\frac{p-1}{2}\leq{dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})<p\end{array}\right.\right..$ (6)
Based on the recovered distances, we denote the global distance matrix as $\boldsymbol{D}\in\mathbb{R}^{n\times n}$, with each entry defined as $\boldsymbol{D}_{ii^{\prime}}={dis}(\boldsymbol{x}_{i},\boldsymbol{x}_{i^{\prime}})$ for $i,i^{\prime}\in[n]$.
Cluster Assignment.
With the recovered global distance matrix $\boldsymbol{D}\in\mathbb{R}^{n\times n}$, the server can directly perform clustering without requiring any modification to existing centralized clustering methods. This characteristic demonstrates the simplicity and flexibility of the proposed OmniFC framework.
Specifically, pairwise-distance-dependent centralized clustering methods—such as spectral clustering (SC) ng2001spectral , DBSCAN ester1996density , hierarchical clustering (HC) gowda1978agglomerative , and k-medoids (KMed) rdusseeun1987clustering —can seamlessly utilize $\boldsymbol{D}$ for model training, owing to their intrinsic reliance on pairwise sample distances during the clustering process. For methods that do not explicitly depend on pairwise relationships—such as k-means (KM) macqueen1967some , fuzzy c-means (FCM) bezdek1984fcm , and nonnegative matrix factorization (NMF) lee2000algorithms —the server employs $\boldsymbol{D}$ as a proxy for the raw features $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ to perform clustering. This allows these algorithms to operate as if on centralized data, while implicitly leveraging the global structure encoded in $\boldsymbol{D}$. These federated extensions of centralized methods built upon OmniFC are denoted as OmniFC-SC, OmniFC-DBSCAN, OmniFC-HC, OmniFC-KMed, OmniFC-KM, OmniFC-FCM, and OmniFC-NMF, respectively. Algorithm 1 in the appendix delineates the pseudocode of OmniFC.
3.3 Privacy Analysis
OmniFC adopts LCC encryption to enhance clustering performance while fortifying data privacy. Although LCC enables clients to obtain global awareness via inter-client sharing of Lagrange-encoded data, it also poses emerging privacy threats, as colluding clients may leverage the shared information to infer others’ private data yu2019lagrange . Hence, evaluating OmniFC’s resistance to client collusion is essential for delineating its practical applicability. Theorem 2 provides a formal guarantee that each data point $\boldsymbol{\tilde{x}}_{i}$ maintains information-theoretic security in the presence of up to $t$ colluding clients, thereby affirming the practical applicability of OmniFC.
Theorem 2.
Given the number of noises $t$, a $t$-private OmniFC is achievable if $m\geq 2l+2t-1$, i.e.,
$I(\boldsymbol{\tilde{x}}_{i};\{\boldsymbol{z}_{i,j}\}_{j\in\boldsymbol{\mathcal{C}}})=0,$ (7)
where $I(\cdot;\cdot)$ denotes the mutual information function, $\boldsymbol{\mathcal{C}}\subset[m]$ and $|\boldsymbol{\mathcal{C}}|\leq t$.
Remark 2.
The condition for achieving $t$-private security in Theorem 2 coincides with that for exact distance reconstruction in Theorem 1, i.e., $m\geq 2l+2t-1$. Consequently, by adhering to this constraint, we can increase the number of noises $t$ to strengthen privacy protection without compromising the precision of distance reconstruction.
4 Experiments
4.1 Experimental Setup
Datasets and Evaluation Criteria.
The proposed OmniFC is assessed using seven benchmark datasets across tabular, visual, temporal, and genomic domains, including Iris fisher1936use , MNIST deng2012mnist , Fashion-MNIST xiao2017fashion , COIL-20 nene1996columbia , COIL-100 nene1996columbia , Pendigits keller2012hics , and 10x_73k zheng2017massively . The chosen datasets encompass diverse modalities, dimensionalities, and cluster patterns, facilitating a comprehensive evaluation of the method’s generalizability in practical scenarios.
The evaluation criteria encompass Normalized Mutual Information (NMI) strehl2002cluster and Kappa liu2019evaluation , with higher scores indicating improved clustering performance. Despite the widespread use of NMI, increasing evidence suggests it may be misleading, whereas Kappa is more reliable liu2019evaluation ; yan2024sda ; yan2025significance . Hence, our analysis is grounded in Kappa-based results, with NMI-based outcomes relegated to the appendix for reference. Details of datasets and evaluation criteria are provided in Appendix C.1.
Baselines.
OmniFC is evaluated in comparison with the federated extensions of several centralized clustering methods, including KM-based (k-FED dennis2021heterogeneity , MUFC pan2023machine ), FCM-based (FFCM stallmann2022towards ), SC-based (FedSC qiao2023federated ), and NMF-based (FedMAvg wang2022federated ) methods. To contextualize the performance of federated clustering against its centralized counterpart, we also present results of vanilla KM, FCM, SC, and NMF under centralized settings, referred to as KM_central, FCM_central, SC_central, and NMF_central, respectively.
Federated Settings.
Following Ref. chung2022federated ; yan2024sda , we simulate diverse federated settings by partitioning the real-world dataset into $k^{\star}$ subsets—each representing a client—and adjusting the non-IID level $p$, where $k^{\star}$ denotes the number of true clusters. Specifically, for each client, a fraction $p$ of its data is sampled from a single cluster, while the remaining $1-p$ portion is drawn uniformly across all clusters. As such, $p=0$ recovers the IID setting, whereas $p=1$ induces a maximally skewed distribution, where each client’s data is fully concentrated within a single cluster. Since OmniFC is immune to the Non-IID degree, the Non-IID level $p$ is indicated solely during comparisons with the existing FC baselines and omitted elsewhere.
Figure 4: Comparison between the ground-truth (top row) and reconstructed (bottom row) pairwise distance matrices. The visual consistency indicates that the proposed OmniFC faithfully recovers the inter-sample similarity.
4.2 Experimental Results
Our experiments center on three key aspects: 1) the comparative advantage of OmniFC over existing approaches; 2) the generality of OmniFC in extending centralized clustering methods; and 3) the sensitivity of OmniFC to hyperparameters. Implementation details are provided in Appendix C.2, and supplementary experimental results are presented in Appendix D.
Table 1: Kappa of clustering methods in different federated scenarios. For each comparison, the best result is highlighted in boldface.
Dataset $\boldsymbol{p}$ SC-based methods KM-based methods FCM-based methods NMF-based methods
SC_central FedSC Ours KM_central k-FED MUFC Ours FCM_central FFCM Ours NMF_central FedMAvg Ours
Iris 0.00 0.95 0.95 0.95 0.95 0.38 0.83 0.95 0.95 0.96 0.95 0.57 0.50 0.95
0.25 0.93 0.95 0.95 0.93 0.95 0.49 0.95 0.50 0.95
0.50 0.85 0.95 0.93 0.79 0.95 0.93 0.95 0.50 0.95
0.75 0.93 0.95 0.95 0.81 0.95 0.96 0.95 0.50 0.95
1.00 0.31 0.95 0.71 0.77 0.95 0.97 0.95 0.50 0.95
MNIST 0.00 0.55 0.53 0.55 0.47 0.43 0.41 0.42 0.50 0.48 0.41 0.46 0.40 0.38
0.25 0.54 0.55 0.45 0.50 0.42 0.52 0.41 0.44 0.38
0.50 0.54 0.55 0.29 0.46 0.42 0.53 0.41 0.39 0.38
0.75 0.58 0.55 0.32 0.47 0.42 0.45 0.41 0.45 0.38
1.00 0.38 0.55 0.47 0.43 0.42 0.48 0.41 0.46 0.38
Fashion-MNIST 0.00 0.53 0.54 0.53 0.50 0.46 0.43 0.51 0.53 0.51 0.50 0.51 0.46 0.49
0.25 0.52 0.53 0.43 0.40 0.51 0.47 0.50 0.46 0.49
0.50 0.54 0.53 0.48 0.50 0.51 0.43 0.50 0.46 0.49
0.75 0.47 0.53 0.45 0.45 0.51 0.50 0.50 0.46 0.49
1.00 0.38 0.53 0.32 0.50 0.51 0.46 0.50 0.46 0.49
COIL-20 0.00 0.61 0.68 0.63 0.64 0.42 0.58 0.64 0.59 0.51 0.59 0.56 0.50 0.61
0.25 0.68 0.63 0.46 0.61 0.64 0.47 0.59 0.51 0.61
0.50 0.73 0.63 0.42 0.57 0.64 0.51 0.59 0.44 0.61
0.75 0.54 0.63 0.41 0.58 0.64 0.55 0.59 0.51 0.61
1.00 0.29 0.63 0.46 0.56 0.64 0.59 0.59 0.52 0.61
COIL-100 0.00 0.54 0.34 0.54 0.49 0.48 0.49 0.56 0.49 0.37 0.53 0.43 0.39 0.50
0.25 0.32 0.54 0.45 0.50 0.56 0.38 0.53 0.38 0.50
0.50 0.32 0.54 0.41 0.48 0.56 0.49 0.53 0.39 0.50
0.75 0.29 0.54 0.41 0.50 0.56 0.48 0.53 0.39 0.50
1.00 0.27 0.54 0.43 0.52 0.56 0.51 0.53 0.38 0.50
Pendigits 0.00 0.72 0.74 0.72 0.61 0.59 0.59 0.62 0.66 0.62 0.66 0.45 0.33 0.72
0.25 0.73 0.72 0.46 0.58 0.62 0.61 0.66 0.33 0.72
0.50 0.72 0.72 0.48 0.60 0.62 0.53 0.66 0.33 0.72
0.75 0.69 0.72 0.33 0.49 0.62 0.49 0.66 0.33 0.72
1.00 0.52 0.72 0.53 0.62 0.62 0.70 0.66 0.33 0.72
10x_73k 0.00 0.89 0.52 0.89 0.85 0.40 0.63 0.56 0.53 0.46 0.55 0.88 0.49 0.82
0.25 0.52 0.89 0.55 0.63 0.56 0.47 0.55 0.49 0.82
0.50 0.52 0.89 0.57 0.62 0.56 0.72 0.55 0.49 0.82
0.75 0.54 0.89 0.37 0.65 0.56 0.64 0.55 0.50 0.82
1.00 0.24 0.89 0.30 0.79 0.56 0.64 0.55 0.50 0.82
count - - 8 27 - 2 9 24 - 13 22 - 5 30
Efficacy Analysis.
To comprehensively validate the efficacy of OmniFC, we simulate five scenarios per dataset: IID ($p=0$), mildly non-IID ($p=0.25$), moderately non-IID ($p=0.5$), highly non-IID ($p=0.75$), and fully non-IID ($p=1$). As shown in Table 1, the proposed OmniFC enables superior federated extensions for both pairwise-distance-dependent SC and methods that do not explicitly depend on pairwise relationships, such as KM, FCM, and NMF. For SC, our extended results attain centralized-level clustering fidelity while remaining robust to diverse Non-IID conditions, owing to lossless pairwise distance reconstruction, which remains unaffected by non-IID severity (see Theorem 1 and Figure 4). For centralized methods not explicitly reliant on pairwise relationships, our extended results generally match—and occasionally exceed—their performance under centralized settings, indicating that the global distance matrix $\boldsymbol{D}$ can serve as an effective surrogate for the raw feature matrix $\boldsymbol{X}$ to perform clustering.
Generality Analysis.
To assess OmniFC’s generalizability in extending centralized clustering methods, we integrate it with three additional methods (KMed, DBSCAN, and HC) that have been well-studied in centralized contexts but remain underexplored in federated settings. Like SC, all three methods perform clustering based on inter-sample pairwise distances. Hence, by utilizing OmniFC’s lossless distance reconstruction, these three methods can be effortlessly integrated into the OmniFC framework to facilitate lossless federated extensions, as shown in Table 2.
Table 2: Kappa of different clustering methods.
Dataset KMed-based methods DBSCAN-based methods HC-based methods
Central Ours Central Ours Central Ours
Iris 0.94 0.94 0.50 0.50 0.94 0.95
MNIST 0.31 0.30 0.21 0.21 0.41 0.41
Fashion-MNIST 0.42 0.42 0.14 0.14 0.45 0.45
COIL-20 0.41 0.42 0.58 0.58 0.45 0.45
COIL-100 0.34 0.34 0.37 0.37 0.48 0.48
Pendigits 0.44 0.45 0.48 0.48 0.57 0.57
10x_73k 0.30 0.30 0.01 0.01 0.28 0.28
Sensitivity Analysis.
To assess the hyperparameter sensitivity of OmniFC, we measure the global distance matrix reconstruction loss—defined as the root-mean-square deviation (RMSE) between the ground-truth and reconstructed pairwise distance matrices—across varying number of clients ($m$), noises ($t$), and segments ($l$). In fact, a theoretical guarantee for this has already been provided in Theorem 1: as long as the condition $m\geq 2l+2t-1$ holds, OmniFC is capable of achieving accurate distance reconstruction. This theoretical result is further substantiated by the empirical evidence presented in Fig. 5.
Figure 5: Hyperparameter sensitivity of the global distance matrix reconstruction loss. The gray-highlighted region denotes hyperparameter settings that violate the condition $m\geq 2l+2t-1$ in Theorem 1, thus precluding distance reconstruction.
5 Conclusion
This work introduces OmniFC, a unified and model-agnostic framework via lossless and secure distance reconstruction. Unlike existing methods that rely on model-specific proxies and suffer from data heterogeneity, OmniFC adopts a distance-based perspective that is decoupled from specific clustering models. Benefit from this, theoretical and empirical results show that this framework improves robustness under non-IID settings and supports the extension of a wide range of centralized clustering algorithms to FC.
Beyond FC, the proposed framework may open broader opportunities across federated learning. In particular, the reconstructed global distance matrix can naturally function as a global affinity graph, offering new possibilities for advancing federated graph learning and other domains where capturing global sample relationships is fundamental.
Appendix A Pseudocode of the Proposed OmniFC
The procedure of OmniFC is formally presented in Algorithm 1. On the client side, each sample $\boldsymbol{\tilde{x}}_{i}$ is independently encoded into $\boldsymbol{z}_{i,j}$ based on Equation (3), and then transmitted to the $j$-th client, where $i\in[n]$ and $j\in[m]$. Then, each client $j$ computes pairwise distances between all encoded representations $\boldsymbol{z}_{i,j}$ and $\boldsymbol{z}_{i^{\prime},j}$ $(i,i^{\prime}\in[n])$ using Equation (4), and transmits the results to the central server. On the server side, the global distance matrix is reconstructed based on Equations (5) and (6), and subsequently utilized by a centralized clustering algorithm to derive the final clustering outcome $\pi^{*}$.
Input: Local datasets $\{\boldsymbol{X}_{j}\}_{j=1}^{m}$, prime number $p$, the number of segments $l$, the number of noises $t$, pre-specified hyperparameters $\{{\alpha}_{o}\}_{o=1}^{l+t}$ and $\{{\beta}_{j}\}_{j=1}^{m}$.
Output: The final partition $\pi^{*}$.
1 Clients execute:
2     Local Lagrange Encoding and Secret Sharing:
3        Each sample $\boldsymbol{\tilde{x}}_{i}$ is encoded via Equation (3), i.e., $\boldsymbol{z}_{i,j}=\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\beta_{j})$, and subsequently
4          transmitted to the $j$-th client, where $i\in[n]$ and $j\in[m]$.
5     Global Distance Reconstruction:
6        Each client $j$ computes pairwise distances between all encoded representations $\boldsymbol{z}_{i,j}$ and
7          $\boldsymbol{z}_{i^{\prime},j}$ $(i,i^{\prime}\in[n])$ using Equation (4), and transmits the results to the central server.
8 Server executes:
9     Global Distance Reconstruction:
10         The server reconstructs the global distance matrix according to Equations (5) and (6).
11
12Cluster assignment:
    The global distance matrix is fed into a centralized clustering method to obtain $\pi^{*}$.
Algorithm 1 OmniFC
Appendix B Proofs of Theorems
Table 3: Notations.
Notation Explanation
$m$ Number of clients.
$\boldsymbol{X}\in\mathbb{R}^{n\times d}$ The centralized dataset $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ consists of $n$ $d$-dimensional samples $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$, which are distributed among $m$ clients, i.e., $\boldsymbol{X}=\bigcup_{j=1}^{m}\boldsymbol{X}_{j}$.
$\boldsymbol{\tilde{X}}\in\mathbb{F}_{p}^{n\times d}$ $\boldsymbol{\tilde{X}}$ denotes the representation of $\boldsymbol{X}$ over the finite field $\mathbb{F}_{p}^{n\times d}$, consisting of $n$ $d$-dimensional samples $\{\boldsymbol{\tilde{x}}_{i}\}_{i=1}^{n}$, where each $\boldsymbol{\tilde{x}}_{i}$ corresponds to the transformed version of $\boldsymbol{x}_{i}$ in $\mathbb{F}_{p}^{d}$.
$l$ Number of data segments.
$t$ Number of noises.
$\boldsymbol{s}_{i,o}\in\mathbb{F}_{p}^{\frac{d}{l}}$ $\boldsymbol{\tilde{x}}_{i}=[\boldsymbol{s}_{i,1}^{T},\boldsymbol{s}_{i,2}^{T},\cdots,\boldsymbol{s}_{i,l}^{T}]^{T}$, where $\boldsymbol{s}_{i,o}$ denotes the $o$-th segment of $\boldsymbol{\tilde{x}}_{i}$ for $o\in[l]=\{1,2,\cdots,l\}$. For $l<o\leq l+t$, $\boldsymbol{s}_{i,o}$ corresponds to the $o$-th noise uniformly sampled from $\mathbb{F}_{p}^{\frac{d}{l}}$.
$\{{\alpha}_{o}\}_{o=1}^{l+t}$ A collection of $l+t$ distinct hyperparameters from $\mathbb{F}_{p}$, predetermined by consensus between all clients and the central server, serves to construct the Lagrange interpolation polynomial.
$\{{\beta}_{j}\}_{j=1}^{m}$ A collection of $m$ distinct hyperparameters from $\mathbb{F}_{p}$, predetermined by consensus between all clients and the central server, serves to encode the local data into $m$ distinct representations.
$\boldsymbol{Z}_{j}\in\mathbb{F}_{p}^{n\times\frac{d}{l}}$ The global encoded dataset possessed by client $j$ $(j\in[m]$). $\boldsymbol{Z}_{j}=[\boldsymbol{z}_{1,j},\boldsymbol{z}_{2,j},\cdots,\boldsymbol{z}_{n,j}]^{T}$, where $\boldsymbol{z}_{i,j}$ $(i\in[n]$) is the encoded representation of $\boldsymbol{\tilde{x}}_{i}$ at client $j$.
Before proving the theorems, we first summarize some notations used throughout the main text and this appendix, and introduce two lemmas from Ref. [37] and [12]. Refer to Table 3 for the notions, with the lemmas delineated below.
Lemma 1.
[37] Given $n$ distinct points $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ with mutually different $x_{i}$, there exists a unique polynomial $\boldsymbol{f}(\boldsymbol{x})$ of degree no greater than $n-1$ that interpolates the data, i.e., $\boldsymbol{f}(\boldsymbol{x}_{i})=\boldsymbol{y}_{i}$.
Lemma 2.
[12] Given the number of noises $t$, and a polynomial $\boldsymbol{f}$ used to compute $\boldsymbol{f}(\boldsymbol{\tilde{X}})$, and the degree of $\boldsymbol{f}$ is denoted as $deg(\boldsymbol{f})$. When $m\geq deg(\boldsymbol{f})(l+t-1)+1$, a $t$-private LCC encryption is achievable, e.g.,
$I(\boldsymbol{\tilde{x}}_{i};\{\boldsymbol{z}_{i,j}\}_{j\in\boldsymbol{\mathcal{C}}})=0,$ (8)
where $I(\cdot;\cdot)$ denotes the mutual information function, $\boldsymbol{\mathcal{C}}\subset[m]$ and $|\boldsymbol{\mathcal{C}}|\leq t$.
Proof of Theorem 1.
With Lemma 1, we prove Theorem 1 as follows.
Proof.
For each distance ${dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j})$ $(j\in[m])$, it can be further formulated as:
${dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j})=\left\|\boldsymbol{z}_{i,j}-\boldsymbol{z}_{i^{\prime},j}\right\|_{2}^{2}=\left\|\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\beta_{j})-\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i^{\prime}}}(\beta_{j})\right\|_{2}^{2},$ (9)
implying that it corresponds to the evaluation of a degree-$2(l+t-1)$ polynomial at $\beta_{j}$. According to Lemma 1, the polynomial can be uniquely interpolated from $12(l+t-1)+$ 1 distinct points. That is, when $m\geq 2l+2t-1$, the polynomial can be interpolated from the set $\{(\beta_{j},{dis}(\boldsymbol{z}_{i,j},\boldsymbol{z}_{i^{\prime},j}))\}_{j=1}^{m}$, and it is exactly $\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\beta)$, i.e.,
$\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\beta)=\left\|\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\beta)-\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i^{\prime}}}(\beta)\right\|_{2}^{2}.$ (10)
Particularly, by assigning $\beta=\alpha_{o}$ $(o\in[l])$, the distance between the $o$-th data segments of $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{\tilde{x}}_{i^{\prime}}$ can be accurately recovered:
$\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\alpha_{o})=\left\|\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i}}(\alpha_{o})-\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i^{\prime}}}(\alpha_{o})\right\|_{2}^{2}=\left\|\boldsymbol{s}_{i,o}-\boldsymbol{s}_{i^{\prime},o}\right\|_{2}^{2}.$ (11)
Consequently, the distance between $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{\tilde{x}}_{i^{\prime}}$ can be precisely reconstructed:
$\sum_{o=1}^{l}\boldsymbol{f}_{\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}}}(\alpha_{o})=\sum_{o=1}^{l}\left\|\boldsymbol{s}_{i,o}-\boldsymbol{s}_{i^{\prime},o}\right\|_{2}^{2}={dis}(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{i^{\prime}})$ (12)
Note that since the above proof does not impose any constraints on the distribution of $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{\tilde{x}}_{i^{\prime}}$ across clients, Equation (12) holds irrespective of how data is distributed among clients. ∎
Proof of Theorem 2.
With Lemma 2, we prove Theorem 2 as follows.
Proof.
We prove Theorem 2 by instantiating Lemma 2 with the specific polynomial structure used in the OmniFC framework.
Recall that Lemma 2 states that a $t$-private LCC encryption is achievable when
$m\geq\deg(\boldsymbol{f})(l+t-1)+1,$
where $\boldsymbol{f}$ is the polynomial used in the encoding scheme, and $l$ is the number of data segments.
In the OmniFC setting, the polynomial $\boldsymbol{f}$ is a quadratic distance-based function of degree 2, i.e., $\deg(\boldsymbol{f})=2$. Plugging this into the general LCC bound yields:
$m\geq 2l+2t-1.$
Therefore, under this condition, the mutual information between any private input $\boldsymbol{\tilde{x}}_{i}$ and the encoded messages observed by up to $t$ colluding clients satisfies:
$I(\boldsymbol{\tilde{x}}_{i};\{\boldsymbol{z}_{i,j}\}_{j\in\boldsymbol{\mathcal{C}}})=0,$
where $\boldsymbol{\mathcal{C}}\subset[m]$ and $|\boldsymbol{\mathcal{C}}|\leq t$.
This guarantees $t$-privacy in the OmniFC framework, thus completing the proof. ∎
Appendix C Experimental Details
All experiments are implemented in Python and executed on a system equipped with an Intel Core i7-12650H CPU, 16GB of RAM, and an NVIDIA GeForce RTX 4060 GPU.
C.1 Datasets and Evaluation Criteria
Table 4: Description of datasets.
Dataset Type Size Image size/Features Class
Iris tabular 150 4 3
MNIST image 70000 $28\times 28$ 10
Fashion-MNIST image 70000 $28\times 28$ 10
COIL-20 image 1440 $128\times 128$ 20
COIL-100 image 7200 $128\times 128$ 100
Pendigits time series 10992 16 10
10x_73k gene 73233 720 8
Datasets.
As shown in Table LABEL:datasets, we select seven benchmark datasets across tabular, visual, temporal, and genomic domains, including Iris [27], MNIST [28], Fashion-MNIST [29], COIL-20 [30], COIL-100 [30], Pendigits [31], and 10x_73k [32]. The chosen datasets encompass diverse modalities, dimensionalities, and cluster patterns, facilitating a comprehensive evaluation of the method’s generalizability in practical scenarios.
Fig. 6 exemplifies, through the Iris dataset, our simulation of federated scenarios under different Non-IID conditions. We simulate diverse federated settings by evenly partitioning the Iris dataset into 3 (the number of true clusters) subsets—each representing a client—and adjusting the non-IID level $p$. For a client with $50$ datapoints, the first $p\cdot 50$ datapoints are sampled from a single cluster, and the remaining $(1-p)\cdot 50$ ones are randomly sampled from any cluster. As such, $p=0$ recovers the IID setting, whereas $p=1$ induces a maximally skewed distribution, where each client’s data is fully concentrated within a single cluster.
Figure 6: Data partition visualization on Iris.
Evaluation Criteria.
Evaluation is based on two metrics—normalized mutual information (NMI) [33] and Kappa [34]—where elevated scores denote superior clustering quality. Despite being widely adopted, NMI has been shown to have limitations, such as the finite size effect, and fails to account for the importance of small clusters [34, 20, 35]. In contrast, Kappa addresses these concerns, making it a more reliable alternative for clustering evaluation. Hence, our analysis is grounded in Kappa-based results, with NMI-based outcomes serving only as supplementary references.
C.2 Implementation Details
All centralized clustering methods are implemented by leveraging existing open-source Python libraries: KM, KMed, SC, NMF, and DBSCAN utilize the sklearn library [38], HC employs the scipy library [39], and FCM adopts an individual open-source implementation [40]. For OmniFC, $\{{\alpha}_{o}\}_{o=1}^{l+t}$ is set as a sequence of $l+t$ consecutive odd integers starting from 1, while $\{{\beta}_{j}\}_{j=1}^{m}$ is set as a sequence of $m$ consecutive even integers starting from 0. The default values of $l$ and $t$ are set to 2.
Additionally, several clustering methods evaluated in our experiments demand full $n\times n$ pairwise distance matrix computations, imposing substantial computational and memory burdens on large-scale datasets. To facilitate the execution of comprehensive experiments, we implement a subsampling strategy whereby 1000 samples are randomly drawn from datasets exceeding 5000 entries to form the experimental subset. This approach balances computational efficiency with the preservation of the original data distribution, enabling fair and meaningful comparisons across methods. The sensitivity of the proposed OmniFC with respect to the number of samples is presented in Appendix D.2.
Table 5: NMI of clustering methods in different federated scenarios. For each comparison, the best result is highlighted in boldface.
Dataset $\boldsymbol{p}$ SC-based methods KM-based methods FCM-based methods NMF-based methods
SC_central FedSC Ours KM_central k-FED MUFC Ours FCM_central FFCM Ours NMF_central FedMAvg Ours
Iris 0.00 0.90 0.90 0.90 0.90 0.66 0.76 0.90 0.90 0.91 0.90 0.56 0.73 0.90
0.25 0.85 0.90 0.90 0.85 0.90 0.72 0.90 0.73 0.90
0.50 0.75 0.90 0.87 0.70 0.90 0.87 0.90 0.73 0.90
0.75 0.85 0.90 0.90 0.74 0.90 0.91 0.90 0.73 0.90
1.00 0.29 0.90 0.70 0.70 0.90 0.93 0.90 0.73 0.90
MNIST 0.00 0.58 0.59 0.58 0.54 0.51 0.48 0.46 0.55 0.53 0.43 0.47 0.48 0.47
0.25 0.60 0.58 0.49 0.52 0.46 0.53 0.43 0.45 0.47
0.50 0.59 0.58 0.39 0.50 0.46 0.52 0.43 0.43 0.47
0.75 0.59 0.58 0.46 0.52 0.46 0.52 0.43 0.47 0.47
1.00 0.45 0.58 0.51 0.55 0.46 0.57 0.43 0.47 0.47
Fashion-MNIST 0.00 0.61 0.61 0.61 0.62 0.56 0.56 0.52 0.61 0.61 0.53 0.60 0.53 0.55
0.25 0.60 0.61 0.54 0.54 0.52 0.59 0.53 0.53 0.55
0.50 0.61 0.61 0.57 0.60 0.52 0.58 0.53 0.53 0.55
0.75 0.55 0.61 0.55 0.54 0.52 0.61 0.53 0.53 0.55
1.00 0.39 0.61 0.48 0.59 0.52 0.58 0.53 0.53 0.55
COIL-20 0.00 0.75 0.80 0.75 0.74 0.65 0.74 0.74 0.75 0.71 0.72 0.70 0.62 0.75
0.25 0.78 0.75 0.70 0.73 0.74 0.69 0.72 0.62 0.75
0.50 0.80 0.75 0.66 0.72 0.74 0.72 0.72 0.62 0.75
0.75 0.69 0.75 0.67 0.73 0.74 0.74 0.72 0.63 0.75
1.00 0.46 0.75 0.69 0.72 0.74 0.75 0.72 0.63 0.75
COIL-100 0.00 0.79 0.67 0.79 0.77 0.76 0.76 0.79 0.79 0.69 0.79 0.72 0.70 0.76
0.25 0.66 0.79 0.75 0.76 0.79 0.71 0.79 0.70 0.76
0.50 0.66 0.79 0.75 0.76 0.79 0.77 0.79 0.70 0.76
0.75 0.64 0.79 0.75 0.76 0.79 0.77 0.79 0.70 0.76
1.00 0.61 0.79 0.75 0.79 0.79 0.81 0.79 0.70 0.76
Pendigits 0.00 0.72 0.77 0.72 0.69 0.67 0.67 0.67 0.69 0.68 0.70 0.55 0.42 0.71
0.25 0.76 0.72 0.62 0.66 0.67 0.68 0.70 0.42 0.71
0.50 0.74 0.72 0.63 0.67 0.67 0.67 0.70 0.42 0.71
0.75 0.75 0.72 0.50 0.64 0.67 0.65 0.70 0.42 0.71
1.00 0.62 0.72 0.64 0.71 0.67 0.69 0.70 0.42 0.71
10X_73k 0.00 0.85 0.71 0.85 0.82 0.68 0.65 0.58 0.68 0.69 0.58 0.83 0.66 0.78
0.25 0.71 0.85 0.70 0.68 0.58 0.70 0.58 0.66 0.78
0.50 0.70 0.85 0.73 0.72 0.58 0.79 0.58 0.66 0.78
0.75 0.59 0.85 0.65 0.73 0.58 0.83 0.58 0.66 0.78
1.00 0.19 0.85 0.49 0.80 0.58 0.82 0.58 0.66 0.78
count - - 13 22 - 5 12 18 - 22 13 - 1 34
Appendix D Supplementary Experimental Results
D.1 NMI-based Evaluation Results
To supplement the Kappa-based evaluation results and to enable broader comparability with existing FC works, we additionally provide NMI-based evaluation results in Tables 5 and 6. Similar to the Kappa-based evaluation results, the numerical results based on NMI also confirm the effectiveness and generalizability of OmniFC.
D.2 Sensitivity Analysis
To assess the sensitivity of the proposed OmniFC concerning the number of samples, we evaluate the global distance matrix reconstruction loss—defined as the root-mean-square deviation (RMSE) between the ground-truth and the reconstructed pairwise distance matrices—across different sample sizes. As shown in Table 7, OmniFC exhibits favorable scalability concerning sample size.
Table 6: NMI of different clustering methods.
Dataset KMed-based methods DBSCAN-based methods HC-based methods
Central Ours Central Ours Central Ours
Iris 0.86 0.86 0.73 0.73 0.89 0.90
MNIST 0.38 0.38 0.56 0.56 0.49 0.49
Fashion-MNIST 0.49 0.49 0.53 0.53 0.54 0.54
COIL-20 0.60 0.61 0.86 0.86 0.70 0.69
COIL-100 0.69 0.69 0.85 0.85 0.78 0.78
Pendigits 0.56 0.55 0.74 0.74 0.69 0.69
10x_73k 0.31 0.31 0.45 0.45 0.51 0.50
Table 7: Sample-size sensitivity of the global distance matrix reconstruction loss on MNIST.
$n$ 1000 2000 3000 5000 7000 10000
RMSE 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
Appendix E Limitation
This work primarily focuses on extending shallow centralized clustering methods and may be less effective for high-dimensional or intrinsically complex data. A promising future direction is to explore how the reconstructed global distance matrix can support the federated extension of deep centralized clustering methods [41], thereby enabling more powerful representation learning under complex data distributions.